# Generalizable Whole Slide Image Classification with Fine-Grained Visual-Semantic Interaction

### 29 Feb 2024 v1

@@ Abstract
Whole Slide Image (WSI) classification is often formulated as a Multiple Instance Learning (MIL) problem. Recently, Vision-Language Models (VLMs) have demonstrated remarkable performance in WSI classification. However, existing methods leverage coarse-grained pathogenetic descriptions for visual representation supervision, which are insufficient to capture the complex visual appearance of pathogenetic images, hindering the generalizability of models on diverse downstream tasks. Additionally, processing high-resolution WSIs can be computationally expensive. In this paper, we propose a novel “Fine-grained Visual-Semantic Interaction” (FiVE) framework for WSI classification. It is designed to enhance the model’s generalizability by leveraging the interplay between localized visual patterns and fine-grained pathological semantics. Specifically, with meticulously designed queries, we start by utilizing a large language model to extract fine-grained pathological descriptions from various non-standardized raw reports. The output descriptions are then reconstructed into fine-grained labels used for training. By introducing a Task-specific Fine-grained Semantics (TFS) module, we enable prompts to capture crucial visual information in WSIs, which enhances representation learning and augments generalization capabilities significantly. Furthermore, given that pathological visual patterns are redundantly distributed across tissue slices, we sample a subset of visual instances during training. Our method demonstrates robust generalizability and strong transferability, dominantly outperforming the counterparts on the TCGA Lung Cancer dataset with at least 9.19% higher accuracy in few-shot experiments.
@@ 1 Introduction
Figure 1: Challenges in WSI-text contrastive learning. Most conventional VLM approaches categorize whole slide images using category-level text descriptions, overlooking intra-class differences, leading to a decline in performance and limitations in generalization capabilities. Instead, we extract fine-grained descriptions from pathological reports as slide-level labels to develop our model, exhibiting detailed variations in each WSI.
@@
Histological Whole Slide Image (WSI) classification plays a crucial role in computational pathology by automating disease diagnosis and subtyping. For high-resolution WSI analysis, Multiple Instance Learning (MIL) has become the dominant method. These methods treat each WSI as a “Bag” sample, and aggregate numerous patches within it as instances for thorough decision-making. Nevertheless, most existing methods [18, 16, 31, 15] only focused on processing image data to conduct WSI classification, neglecting critical pathological insights, particularly the expert textual annotations that accompany these slides.
@@
Recently, Vision-Language Models (VLMs) [25, 11, 33] underscored the significance of integrating multimodal information for developing robust encoders. Zhang et al. [38] exploited disease-grade text labels and extracted text insights using pre-trained language models. Qu et al. [23] utilized GPT-4 in a question-and-answer mode to obtain language prior knowledge at both instance and bag levels for VLM training. However, the challenge lies in the uniqueness and variability of content in each WSI. Existing methods developed their models with coarse-grained descriptions (i.e., simplistic Category-Level text labels [38] or descriptive Category-Level text labels constructed by GPT-4 [23]), as shown in  Fig. 1. They may missed crucial fine-grained pathological details, including differentiation level, vascular invasion, etc., which tends to reduce model performance and limit its generalization.
@@
WSIs accompanied by their corresponding reports (i.e., WSI-report pairs) offer detailed descriptions and fine-grained information vital for WSI analysis. Furthermore, a substantial collection of these pairs is accessible in public databases, such as The Cancer Genome Atlas (TCGA) [9]. Despite this, their full potential has not been adequately harnessed yet. The challenge in developing a VLM using WSI-report pairs mainly lies in the diverse formats and standards of the raw reports from different hospitals, which increases the complexity of data preprocessing and standardization processes. Additionally, pathological reports often contain extraneous information, including report metadata, tissue processing descriptions, and repetitive elements, which introduce noise to textual data. How to extract useful information from raw pathological reports to construct WSI-report pairs is a key issue. In addition, prompt learning has shown effectiveness in improving VLM, which is beneficial in enabling the filtering of image features and successful generalization for downstream tasks. Unlike natural images, WSI data contains a lot of professional pathological information and details. How to craft prompts to make full use of this semantic information to guide fine-grained feature learning is a challenging task. Besides, the high computational cost to train models with high-resolution WSIs also limits the promotion of the model, resulting in a certain resource threshold for WSI analysis.
@@
To address these issues, we proposed a novel whole slide image classification method with Fine-grained Visual-sEmantic interaction termed as FiVE, which shows robust generalizability and efficiency in computation. Firstly, we obtained WSIs with non-standardized raw pathological reports from a public database. Collaborating with professional pathologists, we crafted a set of specialized prompts to standardize reports. Following this, we employed the large language model GPT-4 to automatically clean and standardize the raw report data. In addition, we proposed the Task-specific Fine-grained Semantic (TFS) Module, which utilizes manually designed prompts to direct visual attention to specific pathological areas while constructing Fine-Grained Guidance to enhance the semantic relevance of model features. Considering the diffuse distribution of pathological diseases within tissue sections and the presence of numerous non-diagnostic regions in WSIs, we also incorporated a patch sample strategy during the training phase to enhance training efficiency and reduce computational costs. The contributions of this paper are summarized as follows:
@@
    We introduce a novel Task-specific Fine-grained Semantics (TFS) Module to provide fine-grained guidance for model training, substantially improving the model’s generalization capabilities.
@@ 2.1 Whole Slide Image Analysis
Figure 2: Left: The structure of the FiVE framework. The model consists of a frozen image encoder, a text encoder, and the TFS module. Whole slide images are divided into instances for embedding extraction by the image encoder. Raw pathological reports are standardized by GPT-4 into fine-grained descriptions. The fine-grained descriptions and manual prompts are sampled, shuffled, and reconstructed in pairs. These prompts aggregate instances into bag-level features, subsequently aligned with the descriptions utilizing contrastive loss. Top Right: Fine-grained pathological descriptions. The fine-grained pathological descriptions are generated from multiple answers based on specific queries. These descriptions undergo a process of random sampling, shuffling, and reconstruction to form a unified sentence. Bottom Right: The Instance Aggregator module. The instance aggregator consists of a self-attention module and a cross-attention module, fusing image instance embeddings and prompt embeddings to create bag-level features.
@@ 2.2 Vision-Language Models
Recent research made efforts to develop Vision-Language Models (VLMs). CLIP [25] gathered 400 million image-text pairs and initiated training with synchronized vision and text encoders from the onset. LiT [35] developed a text encoder compatible with a pre-trained vision encoder. FLIP [33] integrated region-word alignment in contrastive learning to enhance detailed vision-language correlation. Coca [34] pre-trained an image-text encoder-decoder foundation model using contrastive and captioning loss. For pathological images, some research works adapted VLMs for training with pathological images and text. Lu et al. [21] built a VLM using over 1.17 million histopathology image-caption pairs based on a task-agnostic pre-training approach derived from Coca. Huang et al. [8] curated the OpenPath dataset, consisting of 208,414 pathology images with natural language descriptions from public forums, and fine-tuned a pre-trained CLIP on OpenPath. Lai et al. [13] also explored the generalization of CLIP in pathology image classification. These methods typically employ instance-level images (small patches from WSIs) and descriptions, requiring significant human and material resources. Zhang et al. [38] injected meaningful medical domain knowledge to advance pathological image embedding and classification. Qu et al. [23] employed GPT-4 to supplement image labels to enrich the information for training. However, the texts employed in their training offer only rudimentary descriptions of the images, primarily categorizing the general type of pathological slides, such as the disease category. This approach significantly constrains the model’s capacity to discern fine-grained features, including the degree of differentiation, spread, and other details. Consequently, this limitation substantially restricts the model’s generalizability and applicability to more nuanced diagnostic tasks.
@@ 2.3 Prompt Learning in Vision-Language Models
Drawing inspiration from prompt learning in natural language processing, some studies have proposed adapting Vision-Language models through end-to-end training of prompt tokens. CoOp [40] enhanced CLIP for few-shot transfer by optimizing a continuous array of prompt vectors within its language branch. Co-CoOp [39] identified CoOp’s suboptimal performance on new classes and tackled the generalization issue by conditioning prompts directly on image instances. Lu et al. [22] advocated for optimizing diverse sets of prompts by understanding their distribution. Bahng et al. [1] undertook visual prompt tuning on CLIP, focusing the prompting on the vision branch. MaPLe [12] investigated the effectiveness of multi-modal prompt learning in order to improve alignment between vision and language representations. Zhang et al. [37] adopted a set of learnable adaption prompts and prepend them to the word tokens at higher transformer layers, efficiently fine-tuning LLaMA with less cost. Furthermore, in the context of WSI classification, prompts function as valuable adjuncts, enriching contextual information and semantic interpretation. The strategic utilization of prompts substantially improved model performance [23].
Manual-Designed Prompts Fine-Grained Text Description Label Example
@@ 6. Are the margins of the excised tissue clear of disease?
Table 1: Manual-Designed prompts and a few examples of resulting fine-grained text description labels. The answers on the right correspond to the prompts on the left. “Unknown” is used as a placeholder when relevant information cannot be found.
@@ 3.1 Overview
Fig. 2 shows the pipeline of our proposed FiVE method. To initiate the process, we collaborated with professional pathologists to establish a set of standards. Following this, we employed GPT-4 to automatically extract and standardize information based on these various standards. During the training phase, we constructed Fine-Grained Guidance by intricately dividing and reconstructing the text description labels and manual prompts in pairs. The combination of manual prompts and learnable prompts forms the Diagnosis Prompts, utilized to enhance the semantic relevance of the features. Subsequently, we utilized the instance aggregator module to fuse instance features with fine-grained prompts, generating Bag-level features, subsequently aligned with the corresponding fine-grained text description labels. Additionally, to reduce computational costs, we implemented the patch sampling strategy, optimizing the model’s training efficiency while minimizing performance loss.
@@ 3.2 Text Standardization via GPT-4
We utilize fine-grained text description labels extracted from pathological reports to align image bag-level features. Though pathological reports are readily accessible from public databases, their content exhibits significant variability depending on the source. Despite these format differences, pathology diagnoses consistently adhere to specific and well-established diagnostic standards. In our work, we developed fine-grained diagnostic criteria under the guidance of professional pathologists to standardize report data and extract fine-grained insights pertinent to pathological diagnosis, aiming to enhance its generalization capabilities substantially.
Our manual-designed prompts aim to extract the morphological characteristics under the microscope, such as the degree of differentiation and lesion invasion, and filter out information irrelevant to the diagnosis. Subsequently, we employed GPT-4 to extract answers from the original diagnosis reports automatically. If the information queried is absent in the pathological reports, “Unknown” is used as the answer. Tab. 1 shows manually designed prompts and two fine-grained text description label examples. Then, we recombine these extracted fine-grained information and integrated it into a complete description of the case image. More details about the prompts used for Text Standardization are provided in Supplementary Material.
@@ 3.3.1 Fine-grained Guidance Construction
We divide the original fine-grained text descriptions into several parts according to manual-designed prompts, followed by random sampling and eliminating “Unknown” tags from the initial labels. Given that the staged diagnostic reports in pathological descriptions are sequence-independent, we shuffled these preliminary labels and reconstructed them into a full-sentence description. During the training, we train the reconstructed text description labels and reconstructed manual prompts in pairs. For example, consider description A: “Lesion differentiation is moderately differentiated; Margins of the excised tissue are clear of disease.” and description B: “Lesion differentiation is moderately differentiated; Margins of the excised tissue are not clear of disease, as the tumor is within the bronchial margin and parenchymal margin.”. When only the first part of each description is sampled, they would be grouped into the same category. However, when sampling the entire sections, they are considered as distinct descriptions.
This strategy offers three key benefits: 1) Effectively alleviating the parent-child relationship in pathology categories. 2) Provide additional hierarchical semantic perspectives to enhance the text encoder’s semantic comprehension ability. 3) Mitigating discrepancies in category annotation due to incomplete diagnostic information.
@@ 3.3.2 Diagnosis Prompts
In addition, manual-designed prompts may have some flaws, potentially failing to comprehensively capture the specific morphological characteristics of the lesion, and the model struggles to generalize towards unseen classes due to the late fusion through the transformer layers. Besides, fine-tuning the model may not always be feasible as it requires training a large number of parameters. Particularly in the case of low-data regimes, where the availability of training data like whole slide images is extremely limited. LLaMA-Adapter [37] and LLaMA-Adapter-v2 [6] explore the way to efficient fine-tuning of Language Models and Vision-Language Models respectively. These approaches introduced the Adaptation Prompt to gradually acquire instructional knowledge. They adopted zero-initialized attention with gating mechanisms to ensure stable training in the early stages. Inspired by these methods, we introduce learnable continuous diagnosis prompts $U_{l}$ to enrich the context information and make our model have stronger transferability.
Specifically, we get the manual text prompt tokens $Q_{h}=[q_{h1},q_{h2},\cdots,q_{hn}]$, here $n$ represent the number of the manul prompts. We concatenate the learnable continuous prompt tokens $Q_{l}=[q_{l1},q_{l2},\cdots,q_{lm}]$ on it, here $m$ represent the number of the learnable prompt tokens. Finally we get the diagnosis prompts $Q=[q_{h1},q_{h2},\cdots,q_{hn},q_{l1},q_{l2},\cdots,q_{lm}]$. In the training phase, part of manual prompt tokens $Q_{h}$ will be sampled randomly, while the whole learnable prompt tokens $Q_{l}$ will be consistently retained for training.
Different with the traditional context learning prompts method, our approach pays attention to the acquisition of prior knowledge, similar to the methodology employed in Detection Transformer (DETR) [4]. We aim to acquire a set of appropriate query values to streamline subsequent feature screening processes, and can also quickly transfer to other tasks by fine-tuning this set of queries.

### 05 Apr 2024 v2

@@ Abstract
Whole Slide Image (WSI) classification is often formulated as a Multiple Instance Learning (MIL) problem. Recently, Vision-Language Models (VLMs) have demonstrated remarkable performance in WSI classification. However, existing methods leverage coarse-grained pathogenetic descriptions for visual representation supervision, which are insufficient to capture the complex visual appearance of pathogenetic images, hindering the generalizability of models on diverse downstream tasks. Additionally, processing high-resolution WSIs can be computationally expensive. In this paper, we propose a novel “Fine-grained Visual-Semantic Interaction" (FiVE) framework for WSI classification. It is designed to enhance the model’s generalizability by leveraging the interaction between localized visual patterns and fine-grained pathological semantics. Specifically, with meticulously designed queries, we start by utilizing a large language model to extract fine-grained pathological descriptions from various non-standardized raw reports. The output descriptions are then reconstructed into fine-grained labels used for training. By introducing a Task-specific Fine-grained Semantics (TFS) module, we enable prompts to capture crucial visual information in WSIs, which enhances representation learning and augments generalization capabilities significantly. Furthermore, given that pathological visual patterns are redundantly distributed across tissue slices, we sample a subset of visual instances during training. Our method demonstrates robust generalizability and strong transferability, dominantly outperforming the counterparts on the TCGA Lung Cancer dataset with at least 9.19% higher accuracy in few-shot experiments. The code is available at: https://github.com/ls1rius/WSI_FiVE.
@@ 1 Introduction
Figure 1: Challenges in WSI-text contrastive learning. Most conventional VLM approaches categorize whole slide images using category-level text descriptions, overlooking intra-class differences, leading to a decline in performance and limitations in generalization capabilities. Instead, we extract fine-grained descriptions from pathology reports as slide-level labels to develop our model, exhibiting detailed variations in each WSI.
@@
Histological Whole Slide Image (WSI) classification plays a crucial role in computational pathology by automating disease diagnosis and subtyping. For high-resolution WSI analysis, Multiple Instance Learning (MIL) has become the dominant method. These methods treat each WSI as a “Bag” sample, and aggregate numerous patches within it as instances for thorough decision-making. Nevertheless, most existing methods [18, 16, 31, 15] focused on processing image data to conduct WSI classification, potentially not emphasizing critical pathological insights, particularly the expert textual annotations that accompany these slides.
@@
Recently, Vision-Language Models (VLMs) [25, 11, 33] underscored the significance of integrating multimodal information for developing robust encoders. Zhang et al. [38] exploited disease-grade text labels and extracted text insights using pre-trained language models. Qu et al. [23] utilized GPT-4 in a question-and-answer mode to obtain language prior knowledge at both instance and bag levels for VLM training. However, the challenge lies in the uniqueness and variability of content in each WSI. Existing methods developed their models with coarse-grained descriptions (i.e., simplistic Category-Level text labels [38] or descriptive Category-Level text labels constructed by GPT-4 [23]), as shown in  Fig. 1. They may have omitted crucial fine-grained pathological details, including differentiation level, vascular invasion, etc., which results in reduced model performance and limited generalization.
@@
WSIs accompanied by their corresponding reports (i.e., WSI-report pairs) offer detailed descriptions and fine-grained information vital for WSI analysis. Furthermore, a substantial collection of these pairs is accessible in public databases, such as The Cancer Genome Atlas (TCGA) [9]. However, their full potential has not been adequately harnessed yet. The challenge in developing a Visual Language Model (VLM) using WSI-report pairs mainly lies in the diverse formats and standards of the raw reports from different hospitals, which increases the complexity of data preprocessing and standardization processes. Additionally, pathology reports often contain extraneous information, including report metadata, tissue processing descriptions, and repetitive elements, which can introduce noise to the textual data. How to extract useful information from raw pathology reports to construct WSI-report pairs is a key issue. Moreover, recent studies have demonstrated the efficacy of prompt engineering in enhancing VLMs. In contrast to natural images, WSI data encompasses extensive professional pathological information and intricate details. How to craft prompts to make full use of this semantic information to guide fine-grained feature learning is a challenging task. Besides, the high computational costs to train models with high-resolution WSIs also limits the promotion of the model, resulting in a certain resource threshold for WSI analysis.
@@
To address these issues, we propose a novel whole slide image classification method with Fine-grained Visual-sEmantic interaction termed as FiVE, which shows robust generalizability and efficiency in computation. Firstly, we obtain WSIs with non-standardized raw pathology reports from a public database. Collaborating with professional pathologists, we craft a set of specialized prompts to standardize reports. Following this, we employ the large language model GPT-4 to automatically clean and standardize the raw report data. In addition, we propose the Task-specific Fine-grained Semantic (TFS) Module, which utilizes manual-designed prompts to direct visual attention to specific pathological areas while constructing Fine-Grained Guidance to enhance the semantic relevance of model features. Considering the diffuse distribution of pathological diseases within tissue sections and the presence of numerous non-diagnostic regions in WSIs, we also incorporate a patch sampling strategy during the training phase to enhance training efficiency and reduce computational costs. The contributions of this paper are summarized as follows:
@@
    We introduce a novel Task-specific Fine-grained Semantics (TFS) Module to offer fine-grained guidance, significantly enhancing the model’s generalization capabilities.
@@ 2.1 Whole Slide Image Analysis
Figure 2: Left: The structure of the FiVE framework. The model consists of a frozen image encoder, a text encoder, and the TFS module. Whole slide images are divided into instances for embedding extraction by the image encoder. Raw pathology reports are standardized by GPT-4 into fine-grained descriptions. The fine-grained descriptions and manual prompts are sampled, shuffled, and reconstructed in pairs. These prompts aggregate instances into bag-level features, subsequently aligned with the descriptions utilizing contrastive loss. Top Right: Fine-grained pathological descriptions. The fine-grained pathological descriptions are generated from multiple answers based on specific queries. These descriptions undergo a process of random sampling, shuffling, and reconstruction to form a unified sentence. Bottom Right: The Instance Aggregator module. The instance aggregator consists of a self-attention module and a cross-attention module, fusing image instance embeddings and prompt embeddings to create bag-level features.
@@ 2.2 Vision-Language Models
Recent researches have made efforts to develop Vision-Language Models (VLMs). CLIP [25] gathered 400 million image-text pairs and initiated training with synchronized vision and text encoders from the onset. LiT [35] developed a text encoder compatible with a pre-trained vision encoder. FLIP [33] integrated region-word alignment in contrastive learning to enhance detailed vision-language correlation. Coca [34] pre-trained an image-text encoder-decoder foundation model using contrastive and captioning loss. For pathological images, some research works adapted VLMs for training with pathological images and text. Lu et al. [21] built a VLM using over 1.17 million histopathology image-caption pairs based on a task-agnostic pre-training approach derived from Coca. Huang et al. [8] curated the OpenPath dataset, consisting of 208,414 pathology images with natural language descriptions from public forums, and fine-tuned a pre-trained CLIP on OpenPath. Lai et al. [13] also explored the generalization of CLIP in pathology image classification. These methods typically employ instance-level images (small patches from WSIs) and descriptions, requiring significant human and material resources. Zhang et al. [38] injected meaningful medical domain knowledge to advance pathological image embedding and classification. Qu et al. [23] employed GPT-4 to supplement image labels to enrich the information for training. However, the texts employed in their training offer only rudimentary descriptions of the images, primarily categorizing the general type of pathological slides, such as the disease category. This approach significantly constrains the model’s capacity to discern fine-grained features, including the degree of differentiation, spread, and other details. Consequently, this limitation substantially restricts the model’s generalizability and applicability to more nuanced diagnostic tasks.
@@ 2.3 Prompt Learning in Vision-Language Models
Drawing inspiration from prompt learning in natural language processing, some studies have proposed adapting Vision-Language models through end-to-end training of prompt tokens. CoOp [40] enhanced CLIP for few-shot transfer by optimizing a continuous array of prompt vectors within its language branch. CoCoOp [39] identified CoOp’s suboptimal performance on new classes and tackled the generalization issue by conditioning prompts directly on image instances. Lu et al. [22] advocated for optimizing diverse sets of prompts by understanding their distribution. Bahng et al. [1] undertook visual prompt tuning on CLIP, focusing the prompting on the vision branch. MaPLe [12] investigated the effectiveness of multi-modal prompt learning in order to improve alignment between vision and language representations. Zhang et al. [37] adopted a set of learnable adaption prompts and prepend them to the word tokens at higher transformer layers, efficiently fine-tuning LLaMA with less cost. Furthermore, in the context of WSI classification, prompts function as valuable adjuncts, enriching contextual information and semantic interpretation. The strategic utilization of prompts substantially improved model performance [23].
Manual-Designed Standards Fine-Grained Text Description Label Examples
@@ 6. Are the margins of the excised tissue clear of disease?
Table 1: Manual-Designed Standards and Fine-Grained Text Description Label Examples. The answers on the right correspond to the standards on the left. “Unknown" is used as a placeholder when relevant information cannot be found.
@@ 3.1 Overview
Fig. 2 shows the pipeline of our proposed FiVE method. To initiate the process, we collaborate with professional pathologists to establish a set of standards. Following this, we employ GPT-4 to automatically extract and standardize information based on these various standards. During the training phase, we construct Fine-Grained Guidance by intricately dividing and reconstructing the text description labels and manual prompts in pairs. The combination of manual prompts and learnable prompts forms the Diagnosis Prompts, which are utilized to enhance the semantic relevance of the features. Subsequently, the instance aggregator module fuses instance features with fine-grained prompts, generating bag-level features, subsequently align with the corresponding fine-grained text description labels. Additionally, to reduce computational costs, we implement the patch sampling strategy, optimizing the model’s training efficiency while minimizing performance loss.
@@ 3.2 Text Standardization via GPT-4
We utilize fine-grained text description labels extracted from pathology reports to align image bag-level features. Though pathology reports are readily accessible from public databases, their content exhibits significant variability depending on the source. Despite these format differences, pathology diagnoses consistently adhere to specific and well-established diagnostic standards. In our work, we develop fine-grained diagnostic criteria under the guidance of professional pathologists to standardize report data and extract fine-grained insights pertinent to pathological diagnosis, aiming to enhance its generalization capabilities substantially.
The manual-designed standards aim to extract the morphological characteristics under the microscope, such as the degree of differentiation and lesion invasion, and filter out information irrelevant to the diagnosis. Subsequently, we employ GPT-4 to automatically extract answers from the original diagnosis reports based on prompts composed of these standards. If the information queried is absent in the pathology reports, “Unknown” is used as the answer. Tab. 1 shows manual-designed standards and two fine-grained text description label examples. Then, we recombine these extracted fine-grained information and integrate it into a complete description of the case image. More details about the prompts used for Text Standardization are provided in Supplementary Material.
@@ 3.3.1 Fine-grained Guidance Construction
During the training process we utilize these manual-designed standards as our manual-designed prompts. We divide the original fine-grained text descriptions into several parts according to manual-designed prompts, followed by random sampling and eliminating “Unknown” tags from the initial labels. Given that the staged diagnostic reports in pathological descriptions are sequence-independent, we shuffle these preliminary labels and reconstruct them into a full-sentence description. During the training, we train reconstructed text description labels and reconstructed manual prompts in pairs. For example, consider description A: “Lesion differentiation is moderately differentiated; Margins of the excised tissue are clear of disease.” and description B: “Lesion differentiation is moderately differentiated; Margins of the excised tissue are not clear of disease, as the tumor is within the bronchial margin and parenchymal margin.”. When only the first part of each description is sampled, they would be grouped into the same category. However, when sampling the entire sections, they are considered as distinct descriptions. Changes in granularity provide diverse perspectives on the visual image, aligning visual image with text descriptions of varying granularities.
This strategy offers three key benefits: 1) Effectively alleviating the parent-child relationship in pathology categories. 2) Providing additional hierarchical semantic perspectives to enhance the text encoder’s semantic comprehension ability. 3) Mitigating discrepancies in category annotation due to incomplete diagnostic information.
@@ 3.3.2 Diagnosis Prompts
In addition, manual-designed prompts may have some flaws, potentially failing to comprehensively capture the specific morphological characteristics of the lesion, and the model struggles to generalize towards unseen classes due to the late fusion through the transformer layers. Besides, fine-tuning the model may not always be feasible as it requires training a large number of parameters. Particularly in the case of low-data regimes, where the availability of training data like whole slide images is extremely limited. LLaMA-Adapter [37] and LLaMA-Adapter-v2 [6] explore the way to efficient fine-tuning of Language Models and Vision-Language Models respectively. These approaches introduced the Adaptation Prompt to gradually acquire instructional knowledge. They adopted zero-initialized attention with gating mechanisms to ensure stable training in the early stages. Inspired by these methods, we introduce learnable continuous diagnosis prompts to enrich the context information and enhance the model’s transferability.
Specifically, we get the manual text prompt tokens $Q_{h}=[q_{h1},q_{h2},\cdots,q_{hn}]$, here $n$ represents the number of the manual prompts. We concatenate the learnable continuous prompt tokens $Q_{l}=[q_{l1},q_{l2},\cdots,q_{lm}]$ on it, here $m$ represents the number of the learnable prompt tokens. Finally we get the diagnosis prompts $Q=[q_{h1},q_{h2},\cdots,q_{hn},q_{l1},q_{l2},\cdots,q_{lm}]$. In the training phase, part of manual prompt tokens $Q_{h}$ will be sampled randomly paired with text description labels, while the whole learnable prompt tokens $Q_{l}$ will be consistently retained.
Different with the traditional context learning prompts method, our approach pays attention to the acquisition of prior knowledge, similar to the methodology employed in Detection Transformer (DETR) [4]. We aim to acquire a set of appropriate query values to improve performance in subsequent feature screening processes. Additionally, it can also enable the model to quickly transfer to other tasks by fine-tuning this set of queries.